A saltwater solution is made by dissolving 3 pounds of salt in 10 gallons of water. At this rate, how many pounds of salt are needed for 25 gallons of water.
Answer by mananth(16946)

let salt required be x
3/10 = x/25
x=3*25/10
x=7.5 lbs. of salt